A ball is released from the top of the tower of height h metres.It takes t seconds to reach the ground.What is the position of the ball in t/3 seconds

Grade:11

1 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Chakrala,

S = u + a(2n-1)/2

h = 10t -5

S1 = 5(2t/3 -1 ) = (h - 10)/3    in t/3 sec.

 

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT BOmbay